# Geometry of convex geometries

### v1

@@ 1 Introduction
Several fundamental combinatorial structures constitute abstract generalizations of geometric settings. Matroids generalize the linear independence in vector spaces, oriented matroids (OMs) capture the combinatorics of regions in a central hyperplane arrangement in ${\mathbb{R}}^{d}$, ample/lopsided sets (AMPs) encode the regions of the arrangement of coordinate hyperplanes intersected with a convex set, and convex geometries/antimatroids represent an abstraction of Euclidean convexity restricted to a finite set. Finally, complexes of oriented matroids (COMs) are a common generalization of oriented matroids and ample sets and capture the combinatorics of regions in an arbitrary hyperplane arrangement in ${\mathbb{R}}^{d}$ restricted to a convex set $K$. Although this geometric model is a desirable property for a respective combinatorial structure, the realizability question is hard. For example, the problem of characterizing which oriented matroids come from hyperplane arrangements is intractable.
@@ 2.1 Set families and systems of sign vectors
The systems of sign vectors generalize set families. Indeed, any set family ($SS(U,\SS)$ can be encoded as a system of sign vectors by setting for each set $SSX\in\SS$, $X_{e}=-1$ if $e\notin X$ and $X_{e}=-1$ if $e\in X$. In this representation, each set $X$ is encoded by a $\{-1,+1\}$-vector. The encoding with $\{-1,0,+1\}$-vectors is useful when encoding cubes of $Q_{n}$. Indeed, each $X$-cube $Q$ with support $Y$ can be encoded by $\{-1,0,+1\}$-vector $X(Q)$, where $X(Q)_{e}=+1$ if $e\in Y$, $X(Q)_{e}=-1$ if $e\in(U\setminus X)\setminus Y$, and $X(Q)_{e}=0$ if $e\in X$.
@@ 2.2 OMs and COMs
We recall the basic theory of OMs and COMs from [5] and [3], respectively. Co-invented by Bland and Las Vergnas [6] and Folkman and Lawrence [14], and further investigated by Edmonds and Mandel [13] and many other authors, oriented matroids represent a unified combinatorial theory of orientations of ordinary matroids, which simultaneously captures the basic properties of sign vectors representing the regions in a hyperplane arrangement in ${\mathbb{R}}^{d}$ and of sign vectors of the circuits in a directed graph. OMs provide a framework for the analysis of combinatorial properties of geometric configurations occurring in discrete geometry and in machine learning. Point and vector configurations, order types, hyperplane and pseudo-line arrangements, convex polytopes, directed graphs, and linear programming find a common generalization in this language. The Topological Representation Theorem of [14] connects the theory of OMs on a deep level to arrangements of pseudohyperplanes and distinguishes it from the theory of ordinary matroids.
Let $\mathcal{L}\subseteq\{-1,0,+1\}^{U}$ be a system of sign vectors. For the sake of this paper we assume that $\mathcal{L}$ is simple, i.e., for all $e\in U$ there exists $X\in\mathcal{L}$ such that $X_{e}\neq 0$. For $X,Y\in\mathcal{L}$, $\operatorname{\mathrm{S}ep}(X,Y)=\{e\in U:X_{e}Y_{e}=-1\}$ is the separator of $X$ and $Y$. The composition of $X$ and $Y$ is the sign vector $X\circ Y$, where for all $e\in U$, $(X\circ Y)_{e}=X_{e}$ if $X_{e}\neq 0$ and $(X\circ Y)_{e}=Y_{e}$ if $X_{e}=0$.
@@
Let $\leq$ be the product ordering on $\{-1,0,+1\}^{U}$ relative to $0\leq-1,+1$. The poset $(\mathcal{L},\leq)$ of an OM $\mathcal{M}=(U,\mathcal{L})$ with an artificial maximum $\hat{1}$ forms the (graded) big face lattice $\mathcal{F}_{\mathrm{big}}(\mathcal{M})$. The topes $\mathcal{T}$ of $\mathcal{L}$ are the co-atoms of $\mathcal{F}_{\mathrm{big}}(\mathcal{M})$. By simplicity the topes are $\{-1,+1\}$-vectors and ${\mathcal{T}}$ can be seen as a family of subsets of $U$. For each $T\in\mathcal{T}$, an element $e\in U$ belongs to the corresponding set if and only if $T_{e}=+1$. The tope graph $G(\mathcal{M})$ of an OM $\mathcal{M}=(U,\mathcal{L})$ is the 1-inclusion graph of the set $\mathcal{T}$ of topes of $\mathcal{L}$, i.e., the subgraph of the hypercube induced by the vertices corresponding to $\mathcal{T}$. Then $G(\mathcal{M})$ is an isometric subgraph of $\mathrm{Q}_{n}$.
@@
One can see that OMs are exactly the COMs containing the zero vector ${\bf 0}$, see [3]. The twist between (Sym) and (FS) allows to keep using the same concepts, such as topes, tope graphs, the sign-order and the big face (semi)lattice in a completely analogous way. On the other hand, it leads to a combinatorial and geometric structure that is build from OMs as cells but is much richer than OMs. Let $\mathcal{M}=(U,\mathcal{L})$ be a COM and $X\in\mathcal{L}$ a covector. The face of $X$ is $\hskip 1.0pt\uparrow X:=\{Y\in\mathcal{L}:X\leq Y\}$, sometimes denoted $\mathrm{F}(X)$, see [3, 5]. A facet is an inclusion maximal proper face. From the definition, any face $\hskip 1.0pt\uparrow X$ consists of the sign vectors of all faces of the subcube of $[-1,+1]^{U}$ with barycenter $X$. By [3, Lemma 4], each face $\hskip 1.0pt\uparrow X$ of a COM $\mathcal{M}$ is an OM. The topes $\mathcal{T}$ and the tope graph $G(\mathcal{M})$ of a COM are defined as for OMs. Again, the COM $\mathcal{M}=(U,\mathcal{L})$ can be recovered from $G(\mathcal{M})$ and $G(\mathcal{M})$ is an isometric subgraph of $\mathrm{Q}_{n}$, see [3, 18].
@@ 2.4 Convex geometries
Convex geometries, introduced and investigated by Edelman and Jamison [11], are the abstract convexity spaces satisfying one of the most important properties of Euclidean convexity: each convex set is the convex hull of its extremal points. A convex geometry [11] is a pair $\mathcal{A}=(U,\mathcal{C})$ of a finite universe $U$ and a collection of convex sets $\mathcal{C}\subseteq 2^{U}$ satisfying the following two conditions:
@@
The convex hull $\operatorname{conv}(A)$ of a set $A\subset U$ is the intersection of all convex sets of $\mathcal{A}$ containing $A$. A point $x$ of a convex set $X$ is called an extreme point of $X$ if $X\setminus\{x\}$ is also convex. Denote by $\operatorname{ex}(X)$ the set of all extreme points of $X$. Under axioms (C1) and (C2), the axiom (C3) is equivalent to the following axiom:
@@
    $X=\operatorname{conv}(\operatorname{ex}(X))$ for any convex set $X$.
Convex geometries can be also characterized by the anti-exchange axiom:
@@
A pair $\mathcal{I}=(U^{\prime},\mathcal{C}^{\prime})$ is called an ideal of a convex geometry if there is a convex geometry $\mathcal{A}=(U,\mathcal{C})$ such that $U^{\prime}\subseteq U$, $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ and if $X\in\mathcal{C},Y\in\mathcal{C}^{\prime}$ and $X\subseteq Y$, then $X\in\mathcal{C}^{\prime}$ (equivalently, for any $X\in\mathcal{C}^{\prime}$, the intervals in $\mathcal{C}^{\prime}$ and $\mathcal{C}$ between $\varnothing$ and $X$ coincide). One can show that ideals of convex geometries are ample sets because they are particular instances of conditional antimatroids, see [2] and Section 6.2. This will also be a consequence of our realizability results.
@@ 3.1 Realizability of COMs
We continue with the definition of realizable COMs given in the paper [3], which generalizes realizability of oriented and affine oriented matroids [5] and of ample sets [22]. An affine arrangement of hyperplanes $\mathcal{H}$ is a finite set of affine hyperplanes of ${\mathbb{R}}^{d}$. We will denote by ${\mathcal{H}}_{d}$ the set of $d$ coordinate hyperplanes of ${\mathbb{R}}^{d}$. Let $K$ be an open convex set of ${\mathbb{R}}^{d}$. Each (oriented) hyperplane $H=\{x\in{\mathbb{R}}^{d}:ax=b\}$ in an arrangement of hyperplanes $\mathcal{H}$ splits $\mathbb{R}^{d}$ into the positive part $H^{+}=\{x\in{\mathbb{R}}^{d}:ax>b\}$, the negative part $H^{-}=\{x\in{\mathbb{R}}^{d}:ax<b\}$, and the zero part $H$. The arrangement $\mathcal{H}$ partitions ${\mathbb{R}}^{d}$ into open convex regions, called cells. All points belonging to the same cell have the same sign vector with respect to the hyperplanes of ${\mathcal{H}}$. Restrict the arrangement pattern to $K$, that is, remove all sign vectors which represent the open regions of the partition disjoint from $K$. The resulting set of sign vectors of the cells of $K$ is denoted $\mathcal{L}({\mathcal{H}},K)$ and constitutes a COM $\mathcal{M}({\mathcal{H}},K)=(H,\mathcal{L}({\mathcal{H}},K))$, see [3].
@@
If ${\mathcal{H}}$ is a central arrangement with $K$ being any open convex set containing the origin, then $\mathcal{M}({\mathcal{H}},K)$ coincides with the notion of realizable oriented matroid [5]. If the arrangement ${\mathcal{H}}$ is affine and $K$ is the entire space, then $\mathcal{M}({\mathcal{H}},K)$ coincides with the realizable affine oriented matroid. Finally, the realizable ample sets arise by taking the central arrangement ${\mathcal{H}}_{d}$ of coordinate hyperplanes restricted to an arbitrary open convex set $K$ of ${\mathbb{R}}^{d}$ (this model was first considered in [22]). A COM $\mathcal{M}$ is called realizable if there exists an affine arrangement of hyperplanes $\mathcal{H}$ and an open convex set $K$ of ${\mathbb{R}}^{d}$ such that $\mathcal{M}=\mathcal{M}({\mathcal{H}},K)$. Then the pair $(\mathcal{H},K)$ is called a realization of $\mathcal{M}$. As noticed in [3], for a realizable COM the open convex set $K$ always can be selected to be an open polyhedron (alias open polyhedral cone). For a realizable COM $\mathcal{M}$, we denote by $\operatorname{dim_{\mathbb{E}}}(\mathcal{M})$ the smallest $d$ such that $\mathcal{M}$ has a realization in $\mathbb{R}^{d}$ and call it the Euclidean dimension of $\mathcal{M}$.
@@
In realizable COMs, $X\leq Y$ for two covectors $X,Y$ if and only if the cell corresponding to $Y$ is contained in the cell corresponding to $X$. Consequently, the topes of realizable COMs are the covectors of the inclusion maximal cells (which all have dimension $d$), called regions. Therefore, the tope graph of a realizable COM with realization $(\mathcal{H},K)$ can be viewed as the adjacency graph of regions: the vertices of this graph are the regions of $K$ defined by the hyperplane arrangement $\mathcal{H}$ and two regions are adjacent in this graph if they are separated by a unique hyperplane of the arrangement (recall that a hyperplane $H$ separates two disjoint open convex sets $A$ and $B$ if $A$ and $B$ belong to distinct halfspaces defined by $H$).
@@
To realize a COM $\mathcal{M}$, it is necessary to find an arrangement of hyperplanes $\mathcal{H}$, an open convex set $K$ (which may be selected to be an open polyhedron), and a sign-preserving bijection between the topes of $\mathcal{M}$ and the sign vectors of the regions (maximal cells) defined by $\mathcal{H}$ and included in $K$. Now, we will show that $\mathcal{H}$ can be selected to be the set of coordinate hyperplanes $\mathcal{H}_{d}$ of ${\mathbb{R}}^{d}$ and $K$ to be an open polyhedron (but not necessarily of full dimension) of ${\mathbb{R}}^{d}$. By this we mean that $K$ is an open set in the induced topology on the affine hull of $K$. This immediately follows from the following lemma:
@@
Let $\mathcal{M}=(U,\mathcal{L})$ be a realizable COM. Then $\mathcal{M}$ has a realization $(\mathcal{H}^{\prime},K^{\prime}),$ where $K^{\prime}$ is an open polyhedron, indeed a cone, and $\mathcal{H}^{\prime}$ is a central arrangement of hyperplanes.
@@
Let $\mathcal{M}=(U,\mathcal{L})$ be a realizable COM with a realization $\mathcal{M}=\mathcal{M}({\mathcal{H}},K)$, where ${\mathcal{H}}$ is an arrangement of affine hyperplanes and $K$ is an open polyhedron of ${\mathbb{R}}^{d}$. First, we consider ${\mathbb{R}}^{d}$ as the hyperplane $x_{d+1}=1$ of ${\mathbb{R}}^{d+1}$. Second, we transform each hyperplane $H_{e}\in{\mathcal{H}},e\in U$ to a central hyperplane $H^{\prime}_{e}$ of ${\mathbb{R}}^{d+1}$ passing via $H_{e}$ and the origin of coordinates. Denote the resulting arrangement by ${\mathcal{H}}^{\prime}$. The supporting hyperplanes $S_{j}$ of the closure of $K$ define a second arrangement of hyperplanes ${\mathcal{S}}$ of ${\mathbb{R}}^{d}$. Analogously to $\mathcal{H}$, the arrangement $\mathcal{S}$ can be transformed into an arrangement ${\mathcal{S}}^{\prime}$ of central hyperplanes of ${\mathbb{R}}^{d+1}$. Then $K$ belongs to one halfspace defined by each hyperplane $S^{\prime}_{j}\in{\mathcal{S}}^{\prime}$. The intersection of those positive open halfspaces is an open polyhedron $K^{\prime}$ containing $K$. Then $\mathcal{L}=\mathcal{L}({\mathcal{H}}^{\prime},K^{\prime})$ holds. Each cell $C\subset K$ of the realization $\mathcal{M}({\mathcal{H}},K)$ gives rise to an open polyhedral cone $C^{\prime}\subset K^{\prime}$ (containing $C$) of the realization $\mathcal{M}({\mathcal{H}}^{\prime},K^{\prime})$. The cone $C^{\prime}$ is the intersection of positive or negative halfspaces defined by the hyperplanes of ${\mathcal{H}}^{\prime}$ and by the halfspaces of the hyperplanes of ${\mathcal{S}}^{\prime}$ containing $K^{\prime}$. ∎
@@
Let $\mathcal{M}=(U,\mathcal{L})$ be a realizable COM. Then $\mathcal{M}$ has a realization $(\mathcal{H}_{n},K^{\prime\prime}),$ where $K^{\prime\prime}$ is an open polyhedron, indeed a cone, and $\mathcal{H}_{n}$ is the set of the coordinates hyperplanes.

### v2

@@ 1 Introduction
Several fundamental combinatorial structures constitute abstract generalizations of geometric settings. Matroids generalize the linear independence in vector spaces, oriented matroids (OMs) capture the combinatorics of regions in a central hyperplane arrangement in ${\mathbb{R}}^{d}$, ample/lopsided sets (AMPs) encode the regions of the arrangement of coordinate hyperplanes intersected with a convex set, and convex geometries/antimatroids represent an abstraction of Euclidean convexity restricted to a finite set. Finally, complexes of oriented matroids (COMs) are a common generalization of oriented matroids and ample sets and capture the combinatorics of regions in an arbitrary hyperplane arrangement in ${\mathbb{R}}^{d}$ restricted to a convex set $K$. Although this geometric model is a desirable property for a respective combinatorial structure, the realizability question is hard. For example, the problem of characterizing which oriented matroids come from hyperplane arrangements is intractable [26, 29, 31] and the realizability of a convex geometry in ${\mathbb{R}}^{2}$ is $\exists{\mathbb{R}}$-complete [2, 18]. All the above structures can be seen as set systems, see Remark 3. This allows to view convex geometries as ample sets and ample sets as COMs.
@@ 2.1 Set families and systems of sign vectors
The systems of sign vectors generalize set families. Indeed, any set family ($SS(U,\SS)$ can be encoded as a system of sign vectors by setting for each set $SSX\in\SS$, $X_{e}=-1$ if $e\notin X$ and $X_{e}=+1$ if $e\in X$. In this representation, each set $X$ is encoded by a $\{-1,+1\}$-vector. The encoding with $\{-1,0,+1\}$-vectors is useful when encoding cubes of $Q_{n}$. Indeed, each $X$-cube $Q$ with support $Y$ can be encoded by $\{-1,0,+1\}$-vector $X(Q)$, where $X(Q)_{e}=+1$ if $e\in Y$, $X(Q)_{e}=-1$ if $e\in(U\setminus X)\setminus Y$, and $X(Q)_{e}=0$ if $e\in X$.
@@ 2.2 OMs and COMs
We recall the basic theory of OMs and COMs from [7] and [5], respectively. Co-invented by Bland and Las Vergnas [8] and Folkman and Lawrence [16], and further investigated by Edmonds and Mandel [15] and many other authors, oriented matroids (OMs) represent a unified combinatorial theory of orientations of ordinary matroids, which simultaneously captures the basic properties of sign vectors representing the regions in a hyperplane arrangement in ${\mathbb{R}}^{d}$ and of sign vectors of the circuits in a directed graph. OMs provide a framework for the analysis of combinatorial properties of geometric configurations occurring in discrete geometry and in machine learning. Point and vector configurations, order types, hyperplane and pseudo-line arrangements, convex polytopes, directed graphs, and linear programming find a common generalization in this language. The Topological Representation Theorem of [16] connects the theory of OMs on a deep level to arrangements of pseudohyperplanes and distinguishes it from the theory of ordinary matroids.
Let $\mathcal{L}\subseteq\{-1,0,+1\}^{U}$ be a system of sign vectors. For the sake of this paper we assume that $\mathcal{L}$ is simple, i.e., $\{X_{e}:X\in\mathcal{L}\}=\{-1,0,+1\}$. For $X,Y\in\mathcal{L}$, $\operatorname{\mathrm{S}ep}(X,Y)=\{e\in U:X_{e}Y_{e}=-1\}$ is the separator of $X$ and $Y$. The composition of $X$ and $Y$ is the sign vector $X\circ Y$, where for all $e\in U$, $(X\circ Y)_{e}=X_{e}$ if $X_{e}\neq 0$ and $(X\circ Y)_{e}=Y_{e}$ if $X_{e}=0$.
@@
One can see that OMs are exactly the COMs containing the zero vector ${\bf 0}$, see [5]. Let $\leq$ be the product ordering on $\{-1,0,+1\}^{U}$ relative to $0\leq-1,+1$. The poset $(\mathcal{L},\leq)$ of a COM $\mathcal{M}=(U,\mathcal{L})$ with an artificial maximum $\hat{1}$ forms the (graded) big face (semi)lattice $\mathcal{F}_{\mathrm{big}}(\mathcal{M})$. For $X\in\mathcal{L}$ a covector of a COM $\mathcal{M}=(U,\mathcal{L})$, the face of $X$ is $\operatorname{\uparrow}X:=\{Y\in\mathcal{L}:X\leq Y\}$, see [5, 7]. By [5, Lemma 4], each face $\operatorname{\uparrow}X$ of a COM $\mathcal{M}$ is an OM.
The topes $\mathcal{T}$ of $\mathcal{L}$ are the co-atoms of $\mathcal{F}_{\mathrm{big}}(\mathcal{M})$. By simplicity the topes are $\{-1,+1\}$-vectors and ${\mathcal{T}}$ can be seen as a family of subsets of $U$. For each $T\in\mathcal{T}$, an element $e\in U$ belongs to the corresponding set if and only if $T_{e}=+1$. Since by [5, 21] the topes determine $\mathcal{M}$ uniquely, we make frequent use of the following:
Remark 3.
Every simple COM $\mathcal{M}$ can be seen as the set family, defined by its set of topes $\mathcal{T}$.
The tope graph $G(\mathcal{M})$ of a COM $\mathcal{M}=(U,\mathcal{L})$ is the 1-inclusion graph of the set $\mathcal{T}$ of topes of $\mathcal{L}$, i.e., the subgraph of the hypercube induced by the vertices corresponding to $\mathcal{T}$. The COM $\mathcal{M}=(U,\mathcal{L})$ can be recovered up to relabelling and reorientation from $G(\mathcal{M})$ and $G(\mathcal{M})$ is an isometric subgraph of $\mathrm{Q}_{n}$, see [5, 21].
@@ 2.4 Convex geometries
Convex geometries, introduced and investigated by Edelman and Jamison [13], are the abstract convexity spaces satisfying one of the most important properties of Euclidean convexity: each convex set is the convex hull of its extremal points. Convex geometries are a particular case of ample sets [4]. A convex geometry [13] is a pair $\mathcal{A}=(U,\mathcal{C})$ of a finite universe $U$ and a collection of convex sets $\mathcal{C}\subseteq 2^{U}$ satisfying the following three conditions:
@@
See the left side of Figure 1 for an example. The convex hull $\operatorname{conv}(A)$ of a set $A\subset U$ is the intersection of all convex sets of $\mathcal{A}$ containing $A$; by (C2), $\operatorname{conv}(A)$ is the smallest element of $\mathcal{C}$ containing $A$. A point $x$ of a convex set $X$ is called an extreme point of $X$ if $X\setminus\{x\}$ is also convex. Denote by $\operatorname{ex}(X)$ the set of all extreme points of $X$. Convex geometries can be also characterized by (C1), (C2), and the following axiom:
@@
    $X=\operatorname{conv}(\operatorname{ex}(X))$ for any $X\in\mathcal{C}$.
Convex geometries can be also characterized by (C1), (C2) and the anti-exchange axiom:
@@
2.5 Bouquets and ideals of convex geometries
We now consider a generalization of ideals of convex geometries, originally introduced under the name of conditional antimatroids in [4]. Analogously to convex geometries bouquets of convex geometries are also ample sets [4]. A pair $\mathcal{A}=(U,\mathcal{C})$ of a finite universe $U$ and a collection of convex sets $\mathcal{C}\subseteq 2^{U}$ is called bouquet of convex geometries if it satisfies the following two conditions:
  • (C1)
    $\varnothing\in\mathcal{C}$;
  • (C2)
    $X\cap Y\in\mathcal{C}$ for all $X,Y\in\mathcal{C}$;
  • (C3)
    for all $X,Y\in\mathcal{C}$ with $Y\subset X$, there is $e\in X\setminus Y$ such that $Y\cup e\in\mathcal{C}$. (locally extendable)
A pair $\mathcal{I}=(U^{\prime},\mathcal{C}^{\prime})$ is called an ideal of a convex geometry if there is a convex geometry $\mathcal{A}=(U,\mathcal{C})$ such that $U^{\prime}\subseteq U$, $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ and if $X\in\mathcal{C},Y\in\mathcal{C}^{\prime}$ and $X\subseteq Y$, then $X\in\mathcal{C}^{\prime}$. Equivalently, for any $X\in\mathcal{C}^{\prime}$, its principal ideals in $\mathcal{C}^{\prime}$ and $\mathcal{C}$ coincide, where the principal ideal $\operatorname{\downarrow}X$ of $X$ in $\mathcal{C}$ (respectively, in $\mathcal{C}^{\prime}$) is the set of all subsets $Y\subseteq X$ such that $Y\in\mathcal{C}$ (respectively, $Y\in\mathcal{C}^{\prime}$). It is easy to see that intervals of bouquets of convex geometries are convex geometries and that ideals of convex geometries are bouquets of convex geometries. However, there are bouquets of convex geometries that are not ideals.
Figure 1: Left: A convex geometry. Right: A bouquet of convex geometries that is not an ideal of a convex geometry. Both are represented in black as a subgraph of the 4-dimensional hypercube (in gray).
Proposition 4.
Proof.
Consider the bouquet of convex geometries in Figure 1, right, i.e., $U^{\prime}=\{1,2,3,4\}$ and
Suppose that $\mathcal{A}^{\prime}=(U^{\prime},\mathcal{C}^{\prime})$ is an ideal of a convex geometry $\mathcal{A}=(U,\mathcal{C})$. Starting with $\{1,4\}$ and applying iteratively (C3), there exists a set $Y\subseteq U\setminus U^{\prime}$ such that one of the sets $Y\cup\{1,2,4\}$ or $Y\cup\{1,3,4\}$ must be in $\mathcal{C}$. Now applying (C2) with respect to $Y\cup\{1,2,4\}$ and $\{2,3,4\}$ or $Y\cup\{1,3,4\}$ and $\{1,2,3\}$, we get that $\{2,4\}$ or $\{1,3\}$ are in $\mathcal{C}$. Hence $\mathcal{A}^{\prime}=(E^{\prime},\mathcal{C}^{\prime})$ is not an ideal of $\mathcal{A}$. ∎
@@ 3.1 Realizability of COMs
We continue with the definition of realizable COMs given in the paper [5], which generalizes realizability of oriented and affine oriented matroids [7] and of ample sets [25]. An affine arrangement of hyperplanes $\mathcal{H}$ is a finite set of affine hyperplanes of ${\mathbb{R}}^{d}$. We will denote by ${\mathcal{H}}_{d}$ the set of $d$ coordinate hyperplanes of ${\mathbb{R}}^{d}$. Let $K$ be a relatively open convex set of ${\mathbb{R}}^{d}$. Each (oriented) hyperplane $H=\{x\in{\mathbb{R}}^{d}:ax=b\}$ in an arrangement of hyperplanes $\mathcal{H}$ splits $\mathbb{R}^{d}$ into the positive part $H^{+}=\{x\in{\mathbb{R}}^{d}:ax>b\}$, the negative part $H^{-}=\{x\in{\mathbb{R}}^{d}:ax<b\}$, and the zero part $H$. The arrangement $\mathcal{H}$ partitions ${\mathbb{R}}^{d}$ into relatively open convex regions, called cells. All points belonging to the same cell have the same sign vector with respect to the hyperplanes of ${\mathcal{H}}$. Restrict the arrangement pattern to $K$, that is, remove all sign vectors which represent the regions of the partition disjoint from $K$. The resulting set of sign vectors of the cells of $K$ is denoted $\mathcal{L}({\mathcal{H}},K)$ and constitutes a COM $\mathcal{M}({\mathcal{H}},K)=(H,\mathcal{L}({\mathcal{H}},K))$, see [5].
@@
If ${\mathcal{H}}$ is a central arrangement with $K$ being any relatively open convex set containing the origin, then $\mathcal{M}({\mathcal{H}},K)$ coincides with the notion of realizable oriented matroid [7]. If the arrangement ${\mathcal{H}}$ is affine and $K$ is the entire space, then $\mathcal{M}({\mathcal{H}},K)$ coincides with the realizable affine oriented matroid. Finally, the realizable ample sets arise by taking the central arrangement ${\mathcal{H}}_{d}$ of coordinate hyperplanes restricted to a full-dimensional open convex set $K$ of ${\mathbb{R}}^{d}$ (this model was first considered in [25]). A COM $\mathcal{M}$ is called realizable if there exists an affine arrangement of hyperplanes $\mathcal{H}$ and a relatively open convex set $K$ of ${\mathbb{R}}^{d}$ such that $\mathcal{M}=\mathcal{M}({\mathcal{H}},K)$. Then the pair $(\mathcal{H},K)$ is called a realization of $\mathcal{M}$. As noticed in [5], for a realizable COM the set $K$ always can be selected to be a relatively open polyhedron. For a realizable COM $\mathcal{M}$, we denote by $\operatorname{dim_{\mathbb{E}}}(\mathcal{M})$ the smallest $d$ such that $\mathcal{M}$ has a realization in $\mathbb{R}^{d}$ and call it the Euclidean dimension of $\mathcal{M}$.
@@
In realizable COMs, $X\leq Y$ for two covectors $X,Y$ if and only if the cell corresponding to $Y$ is contained in the cell corresponding to $X$. Consequently, the topes of realizable COMs are the covectors of the inclusion maximal cells (which all have dimension $d$), called regions. Therefore, the tope graph of a realizable COM with realization $(\mathcal{H},K)$ can be viewed as the adjacency graph of regions: the vertices of this graph are the regions of $K$ defined by the hyperplane arrangement $\mathcal{H}$ and two regions are adjacent in this graph if they are separated by a unique hyperplane of the arrangement (recall that a hyperplane $H$ separates two disjoint relatively open convex sets $A$ and $B$ if $A$ and $B$ belong to distinct halfspaces defined by $H$).
@@
To realize a COM $\mathcal{M}$, it is necessary to find an arrangement of hyperplanes $\mathcal{H}$, a relatively open convex set $K$, and a sign-preserving bijection between the topes of $\mathcal{M}$ and the sign vectors of the regions (maximal cells) defined by $\mathcal{H}$ and included in $K$. Now, we will show that $\mathcal{H}$ can be selected to be the set of coordinate hyperplanes $\mathcal{H}_{d}$ of ${\mathbb{R}}^{d}$ and $K$ to be a relatively open polyhedron of ${\mathbb{R}}^{d}$. This immediately follows from the following lemma:
@@
Let $\mathcal{M}=(U,\mathcal{L})$ be a realizable COM. Then $\mathcal{M}$ has a realization (′′$(\mathcal{H}^{\prime\prime},K^{\prime\prime}),$ where $K^{\prime\prime}$ is a relatively open polyhedral cone, and $\mathcal{H}^{\prime\prime}$ is a central arrangement of hyperplanes, whose normal vectors generate the ambient space.
@@
Let $\mathcal{M}=(U,\mathcal{L})$ be a realizable COM with a realization $\mathcal{M}=\mathcal{M}({\mathcal{H}},K)$, where ${\mathcal{H}}$ is an arrangement of affine hyperplanes and $K$ is a relatively open polyhedron of ${\mathbb{R}}^{d}$, as provided by [5]. First, we consider ${\mathbb{R}}^{d}$ as the hyperplane $x_{d+1}=1$ of ${\mathbb{R}}^{d+1}$. Second, we transform each hyperplane $H_{e}\in{\mathcal{H}},e\in U$ to a central hyperplane $H^{\prime}_{e}$ of ${\mathbb{R}}^{d+1}$ passing via $H_{e}$ and the origin of coordinates. Denote the resulting arrangement by ${\mathcal{H}}^{\prime}$. The supporting hyperplanes $S_{j}$ of the closure of $K$ define a second arrangement of hyperplanes ${\mathcal{S}}$ of ${\mathbb{R}}^{d}$. Analogously to $\mathcal{H}$, the arrangement $\mathcal{S}$ can be transformed into an arrangement ${\mathcal{S}}^{\prime}$ of central hyperplanes of ${\mathbb{R}}^{d+1}$. Then $K$ belongs to one halfspace defined by each hyperplane $S^{\prime}_{j}\in{\mathcal{S}}^{\prime}$. The intersection of those positive open halfspaces is an open polyhedron $K^{\prime}$ containing $K$. Then $\mathcal{L}=\mathcal{L}({\mathcal{H}}^{\prime},K^{\prime})$ holds. Each cell $C\subset K$ of the realization $\mathcal{M}({\mathcal{H}},K)$ gives rise to an open polyhedral cone $C^{\prime}\subset K^{\prime}$ (containing $C$) of the realization $\mathcal{M}({\mathcal{H}}^{\prime},K^{\prime})$. The cone $C^{\prime}$ is the intersection of positive or negative halfspaces defined by the hyperplanes of ${\mathcal{H}}^{\prime}$ and by the halfspaces of the hyperplanes of ${\mathcal{S}}^{\prime}$ containing $K^{\prime}$.
If the normal vectors of ${\mathcal{H}}^{\prime}$ generate a proper subspace $W\subset\mathbb{R}^{d+1}$ of dimension $d^{\prime}$, then we consider ${\mathcal{H}}^{\prime}$ as an arrangement in $W$, intersect $K^{\prime}$ with $W$, and affinely transform $W$ to $\mathbb{R}^{d^{\prime}}$. This yields a realization (′′$(\mathcal{H}^{\prime\prime},K^{\prime\prime})$ of $\mathcal{M}$ with the desired properties. ∎
@@
Let $\mathcal{M}=(U,\mathcal{L})$ be a realizable COM. Then $\mathcal{M}$ has a realization $(\mathcal{H}_{n},K),$ where $K$ is a relatively open polyhedral cone, and $\mathcal{H}_{n}$ is the set of the coordinates hyperplanes.